# Part II: Variational bootstrapping of the $\Lambda$CDM model

### v1

@@ IV Seed for modified field equations
We will now discuss the construction of an “educated guess” for a minimal metric-affine extension of the dynamics of the $\Lambda$CDM model. Therefore, for the rest of this paper, we consider the geometric setup as given by a triple $\left(M,\mathbf{L},\boldsymbol{g}\right)$ as presented in Section II.
In any metric-affine theory of gravity employing the metric and distortion tensor components as dynamical variables, the evolution equations for the metric can be cast in the form
@@
is a symmetric tensor field of type $(0,2)$ (or $(2,0)$) depending on the metric $\boldsymbol{g}$, the distortion tensor components and a finite number of derivatives thereof.
In the particular case of cosmological symmetry, we want $\mathbf{\Theta}^{\text{(D)}}$ to geometrically encode the dark sector contributions, in other words, to be the geometric counterpart of $\boldsymbol{T}^{\text{(D)}}$. To stay as close as possible to the $\Lambda$CDM model, we may want to regard dark matter and dark energy as distinct effects. [^2] This point of view is motivated by the fact that effects attributed to dark-matter are already present at scales smaller than the cosmological one; typically the necessity to introduce dark matter in our physical theory first appeared in the context of rotation curve of galaxies, independently of the presence or not of a cosmological constant (dark-energy) [13]. This corresponds to the idea of writting (31) with
$\mathbf{\Theta}^{\text{(D)}}\coloneqq\mathbf{\Theta}^{\text{(DM)}}+\mathbf{\Theta}^{\text{(DE)}}.$ (33)
In this case, the success of the $\Lambda$CDM model should be seen as constraining the form of $\mathbf{\Theta}^{\text{(DM)}}$ and $\mathbf{\Theta}^{\text{(DE)}}$. In other words, one should make sure that, under cosmological symmetry, the form of $\mathbf{\Theta}^{\text{(DM)}}$ and $\mathbf{\Theta}^{\text{(DE)}}$ are compatible with the dark-matter and dark-energy EOS from the $\Lambda$CDM model
@@
As already stated, our aim in this paper is to consider a minimal metric-affine extension of the $\Lambda$CDM model. This corresponds to the idea that the degrees of freedom describing dark matter and dark energy in a cosmologically symmetric situation are the only ones necessary to desribe the dynamics of these entities in all situations.
In other words, stepping now aside from the cosmological symmetry context, under this “minimality” hypothesis, no extra quantities or terms appear, hence we can simply replace $\overset{\bullet}{=}$ by a strict equality. This gives
@@
    The above postulated equations are not necessarily variational – i.e., they generally do not arise from an action principle.
@@ V Variational Bootstrapping of $\Lambda$CDM
The variational bootstrapping method [12] allows one to determine a Lagrangian whose Euler-Lagrange equations are the closest to an “educated guess” for a subset of the field equations. After briefly reviewing the algorithm, we apply it to (39).
@@ V.1 User’s guide on Variational Bootstrapping
Let us start by presenting a minimalist guide for the method of variational bootstrapping, focussing on computational aspects. For complete and mathematically rigorous details, we refer the reader to [12].
@@
Assume now that we have some insight (an “educated guess”) regarding the form of the $y^{A}$-equation only and that this insight is of the form
@@
In cases where our guessed equations (41) can be obtained from a variational principle, any Lagrangian density producing (41) as $y^{A}$-part of the Euler-Lagrange equations must be of the form (42) up to Lagrangian densities completely independent of the $y^{A}$-variables (i.e. which do not contribute to the $y^{A}$-equations) and total divergences (which do not contribute to the field equations at all), see [12]. The form of the $y^{A}$-independent densities is not constrained by the above procedure and should thus be found by different means [^4] Intuitively, the situation is similar to try finding a function $f:\mathbb{R}^{2}\to\mathbb{R}^{2}:(x,y)\mapsto f(x,y)$ when the only information we have is on $\partial_{x}f$..
But what if the equation (41) we start with is not variational (i.e. what if it does not appear as the $y^{A}$-equation of any Lagrangian) ? In that case, the idea behind variational bootstrapping is to still compute the Lagrangian density (42). This does two things for us:
@@
    It provides a Lagrangian density that will “correct the $y^{A}$-equations”555i.e. the $y^{A}$-part of the Euler-Lagrange equations associated with (42) will be as close as possible from (41), while being variational; the difference between (41) and the $y^{A}$-part of the Euler-Lagrange equations associated with (42) being the components of the so-called Helmholtz form which, in a nutshell, precisely measures the obstruction from variationality of (41).,
@@
In the case of modified gravity theories based on GR, the field variables are usually the metric $\boldsymbol{g}$ and, generally speaking, another set of tensorial field variables, say $\boldsymbol{\psi}^{A}$. Our knowledges on metric-based approaches to gravity provides us with an “educated guess” for the metric equation but not necessarilly for the dynamics of the $\boldsymbol{\psi}^{A}$ fields. [^6] One might have an idea of the form of the dynamical equation for the $\boldsymbol{\psi}^{A}$ variables if those fields have some known dynamics e.g. on Minkowski spacetime which we try to extend in the presence of a generic metric (think e.g. to a scalar field). In such a case, we verified that the variational bootstrapping procedure was in line with the usual approaches, see [12]. One can then to apply variational bootstrapping with respect to the metric to obtain a Lagrangian density based on this guessed equation and then deduce the $\boldsymbol{\psi}^{A}$ dynamics from it.
@@
where $T^{\mu\nu}\left(\boldsymbol{g},\partial\boldsymbol{g};\boldsymbol{\psi}^{A},\partial\boldsymbol{\psi}^{A}\right)$ stands for the energy-momentum tensor associated with the extra fields $\boldsymbol{\psi}^{A}$, the procedure prescribes to compute777where the lower integration bound is here understood as a limit, see [12].
@@
Here again, this procedure provides the “best” Lagrangian density related to our initial guess (43) up to terms that would not give any contribution to the metric equation; i.e. Lagrangian densities built without using the metric $\boldsymbol{g}$. Such Lagrangian densities are quite uncommon for physical theories as one usually uses the metric in the construction of the Lagrangian either to manipulate indices or to define a volume element which is used in the construction of the action. They are, nevertheless, not ruled out a priori by the above construction. In [12], we classified all possible natural Lagrangians independent of the metric in the context of a metric-affine setup in $4$ dimensions. It turns out there is a limited amount of these which are obtained from specific $4$-forms constructed solely from the components of the distorsion tensor $\mathbf{L}$ and their first derivatives, see [12] for a complete classification.
@@ V.2 General requirements
    The quantities $\Lambda$ and $V$ must depend on $\boldsymbol{g}$, $\mathbf{L}$, and their first order derivatives $\partial\boldsymbol{g}$, $\partial\mathbf{L}$ only and in a generaly covariant way, more precisely:  [^8] Note that this hypothesis does not contradict the fact that our expression could depend on $\nabla\mathbf{L}$ since $\nabla_{\alpha}L^{\rho}_{\phantom{\rho}\mu\nu}=\accentset{\circ}{\nabla}_{\alpha}L^{\rho}_{\phantom{\rho}\mu\nu}+L^{\rho}_{\phantom{\rho}\beta\alpha}L^{\beta}_{\phantom{\beta}\mu\nu}-L^{\beta}_{\phantom{\beta}\mu\alpha}L^{\rho}_{\phantom{\rho}\beta\nu}-L^{\beta}_{\phantom{\beta}\nu\alpha}L^{\rho}_{\phantom{\rho}\mu\beta}.$ We can then write $\Lambda=\Lambda(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})$ without loss of generality.
@@
    were each $\Lambda_{(k)}$ is a homogeneous function of degree $k$ in $\boldsymbol{g}$ and its derivative; and the same property holds for $V$.
@@ V.3 Bootstrapping of $\Lambda$CDM
We want to construct here the Lagrangian density obtained by variational bootstrapping of (39). According to (44) and the above hypotheses, we thus need to compute
@@
Here again, according to hypothesis 3, we want to parametrize $V^{\mu}(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})$ as
@@ V.4.1 Variation with respect to $\boldsymbol{g}$
where $\simeq$ means equality up to boundary terms.
Let us now come to the variation of (54) with respect to the metric. To this purpose, it is sufficient to consider
@@
where, to shorten the notation, we have defined $\left(\partial\Lambda_{(k)}\right)^{\alpha\phantom{\rho}\sigma\beta}_{\phantom{\alpha}\rho}\coloneqq\frac{\partial\Lambda_{(k)}}{\partial\left(\accentset{\circ}{\nabla}_{\alpha}L^{\rho}_{\phantom{\rho}\sigma\beta}\right)}$. Using the relations from Appendix B (in particular (93) and (95)), we get that
$\delta_{g}\left(\Lambda_{(k)}\sqrt{-\lvert g\rvert}\right)\simeq\sqrt{-\lvert g\rvert}\ T^{(\mu\nu)}_{(k)}\ \delta g_{\mu\nu},$ (61)
where
$\begin{split}T^{\mu\nu}_{(k)}\coloneqq&\ \frac{1}{2}\Lambda_{(k)}g^{\mu\nu}+\frac{\partial\Lambda_{(k)}}{\partial g_{\mu\nu}}-\accentset{\circ}{\nabla}_{\alpha}\left[L^{\lceil\nu\rvert}_{\phantom{\lceil\nu\rvert}\sigma\beta}\ \left(\partial\Lambda_{(k)}\right)^{\lvert\alpha\rvert\phantom{\rho}\sigma\beta}_{\phantom{\lvert\alpha\rvert}\rho}\ g^{\rho\lvert\mu\rfloor}\right.\\ &\left.-\left\{\left(\partial\Lambda_{(k)}\right)^{\lceil\nu\rvert\phantom{\rho}\lvert\alpha\rvert\beta}_{\phantom{\lceil\nu\rvert}\rho}\ g^{\sigma\lvert\mu\rfloor}+\left(\partial\Lambda_{(k)}\right)^{\lceil\nu|\phantom{\rho}\sigma\lvert\alpha\rvert}_{\phantom{\lceil\nu|}\rho}\ g^{\beta\lvert\mu\rfloor}\right\}L^{\rho}_{\phantom{\rho}\sigma\beta}\right],\end{split}$ (62)
and, for a generic tensor field with components $S^{\rho\mu\nu}$, we have denoted:
$S^{\left\lceil\rho\mu\nu\right\rfloor}\coloneqq\frac{1}{2}\left(S^{\rho\mu\nu}-S^{\nu\rho\mu}+S^{\mu\nu\rho}\right).$ (63)
@@
$\delta_{g}\left(V_{(kl)}\sqrt{-\lvert g\rvert}\right)\simeq\sqrt{-\lvert g\rvert}\ T^{(\mu\nu)}_{(kl)}\ \delta g_{\mu\nu},$ (65)
where the expression for $T^{\mu\nu}_{(kl)}$ is obtained by performing the replacement $\Lambda_{(k)}\to V_{(kl)}$ in (62).
Combining (59), (61) and (65), we get that the metric field equation for lagrangian (46), in the absence of matter source, is
@@
where $\boldsymbol{T}^{(k)}=T^{(k)}_{(\mu\nu)}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ and $\boldsymbol{T}^{(kl)}=T^{(kl)}_{(\mu\nu)}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$.
@@ V.4.2 Variation with respect to $\mathbf{L}$
$\delta_{L}\left(\Lambda_{(k)}\sqrt{-\lvert g\rvert}\right)=\delta_{L}\Lambda_{(k)}\ \sqrt{-\lvert g\rvert}\simeq\sqrt{-\lvert g\rvert}\ \left(\Psi_{(k)}\right)^{\phantom{\rho}\sigma\beta}_{\rho}\ \delta L^{\rho}_{\phantom{\rho}\sigma\beta},$ (67)

### v2

@@ IV Seed for modified field equations
We will now discuss the construction of an “educated guess” for a minimal metric-affine extension of the dynamics of the $\Lambda$CDM model, according to the principles stated in the Introduction. Therefore, for the rest of this paper, we consider the geometric setup as given by a triple $\left(M,\mathbf{L},\boldsymbol{g}\right)$ as in Section II.
We start by noting that, in any metric-affine theory of gravity employing the metric and distortion tensor components as dynamical variables, the evolution equations for the metric can be cast into the form
@@
is a symmetric tensor field of type $(0,2)$ (or $(2,0)$) depending on the metric $\boldsymbol{g}$, the distortion tensor components and a finite number of derivatives thereof. Our work will be to determine an “educated guess” for $\mathbf{\Theta}^{\text{(D)}}$.
Step 1. In the particular case of cosmological symmetry, we want $\mathbf{\Theta}^{\text{(D)}}$ to geometrically encode the dark sector contributions, in other words, to be the geometric counterpart of $\boldsymbol{T}^{\text{(D)}}$. As already discussed above, in cosmological symmetry, the most general possible form of $\mathbf{\Theta}^{\text{(D)}}$ is
$\mathbf{\Theta}^{\text{(D)}}\coloneqq\mathbf{\Theta}^{\text{(DM)}}+\mathbf{\Theta}^{\text{(DE)}},$ (33)
where $\mathbf{\Theta}^{\text{(DM)}}$ and $\mathbf{\Theta}^{\text{(DE)}}$ are interpreted as geometric “dark matter” and “dark energy”. [^3] This point of view is motivated by the fact that effects attributed to dark-matter are already present at scales smaller than the cosmological one; typically the necessity to introduce dark matter in our physical theory first appeared in the context of rotation curve of galaxies, independently of the presence or not of a cosmological constant (dark-energy) [13]. It is then sound to still model these as distinct physical effects. In other words, the following equations of state should hold
@@
Step 2. Leaving now the cosmologically symmetric context and going back to the general one, we build a minimal metric-affine extension of the $\Lambda$CDM model equations, in the sense that the terms (“degrees of freedom”) describing dark matter and dark energy in a cosmologically symmetric situation are the only ones necessary to desribe the dynamics of these entities in all situations. Under this “minimality” hypothesis, no extra quantities or terms appear, hence we can simply replace $\overset{\bullet}{=}$ by a strict equality. This gives
@@
    The above postulated equations are not necessarily variational – that is, they generally do not arise from an action principle.
@@ V Variational Bootstrapping of $\Lambda$CDM
The variational bootstrapping method [12] allows one to determine a Lagrangian whose Euler-Lagrange equations are the closest to an “educated guess” for a given subset of the field equations. After briefly reviewing the algorithm, we apply it to (39).
@@ V.1 User’s guide on Variational Bootstrapping
Below, we present a minimalist guide for the method of variational bootstrapping, focusing on computational aspects. For complete and mathematically rigorous details, we refer the reader to [12].
@@
Assume now that we have some insight (an “educated guess”) regarding the form of the $y^{A}$-equations only and that this insight is of the form
@@
In cases where our guessed equations (41) can be obtained from a variational principle, any Lagrangian density producing (41) as $y^{A}$-part of the Euler-Lagrange equations must be of the form (42) up to Lagrangian densities completely independent of the $y^{A}$-variables (i.e. which do not contribute to the $y^{A}$-equations) and total divergences (which do not contribute to the field equations at all), see [12]. The form of the $y^{A}$-independent terms in the Lagrangian density is not constrained by the above procedure and should thus be found by different means [^5] Intuitively, the situation is similar to try finding a function $f:\mathbb{R}^{2}\to\mathbb{R}^{2}:(x,y)\mapsto f(x,y)$ when the only information we have is on $\partial_{x}f$..
But what if the equation (41) we start with is not variational (i.e. what if it does not arise as the Euler-Lagrange equation with respect to $y^{A}$ of any Lagrangian)? In that case, the idea behind variational bootstrapping is to still compute the Lagrangian density (42). This does two things for us:
@@
    It provides a Lagrangian density that will “correct the $y^{A}$-equations” into their “closest” variational equations. [^6] To be more precise, the difference between (41) and the $y^{A}$-part of the Euler-Lagrange equations associated with (42) is a linear combination of the components of the so-called Helmholtz form, which precisely measures the obstruction from variationality of (41).,
@@
In the case of modified gravity theories based on GR, the field variables are usually the metric $\boldsymbol{g}$ and, generally speaking, another set of tensor variables, say $\boldsymbol{\psi}^{A}$. Our knowledge on metric-based approaches to gravity provides us with an “educated guess” for the metric equation, but not necessarily for the dynamics of the $\boldsymbol{\psi}^{A}$ fields. [^7] One might have an idea of the form of the dynamical equation for the $\boldsymbol{\psi}^{A}$ variables if those fields have some known dynamics (e.g., on Minkowski spacetime), which we try to extend in the presence of a generic metric (think e.g. to a scalar field). In such a case, we verified that the variational bootstrapping procedure was in line with the usual approaches, see [12]. One can then apply variational bootstrapping with respect to the metric to obtain a Lagrangian density based on this guessed equation and then deduce the $\boldsymbol{\psi}^{A}$ dynamics from it.
@@
where $T^{\mu\nu}\left(\boldsymbol{g},\partial\boldsymbol{g};\boldsymbol{\psi}^{A},\partial\boldsymbol{\psi}^{A}\right)$ stands for the energy-momentum tensor associated with the extra fields $\boldsymbol{\psi}^{A}$, the procedure prescribes to compute888where the lower integration bound is here understood as a limit, see [12].
@@
Here again, this procedure provides (up to boundary terms) the “best” Lagrangian density related to our initial guess (43), up to terms that would not give any contribution to the metric equation – that is, Lagrangian densities built without using the metric $\boldsymbol{g}$ at all. Such Lagrangian densities are quite uncommon for physical theories – as one usually uses the metric in the construction of the Lagrangian density to manipulate indices and/or to define a volume element. They are, nevertheless, not ruled out a priori by the above construction. In [12], we classified all possible 4-dimensional natural Lagrangians that are independent of the metric in the context of a metric-affine setup. It turns out there is a limited amount of these which are obtained from specific $4$-forms constructed solely from the components of the distorsion tensor $\mathbf{L}$ and their first derivatives, see [12] for a complete classification.
@@ V.2 General requirements
    The quantities $\Lambda$ and $V$ must depend on $\boldsymbol{g}$, $\mathbf{L}$, and their first order derivatives $\partial\boldsymbol{g}$, $\partial\mathbf{L}$ only and in a generaly covariant way, more precisely: [^9] This hypothesis does not contradict the fact that our expression could depend on $\nabla\mathbf{L}$, since $\nabla_{\alpha}L^{\rho}_{\phantom{\rho}\mu\nu}=\accentset{\circ}{\nabla}_{\alpha}L^{\rho}_{\phantom{\rho}\mu\nu}+L^{\rho}_{\phantom{\rho}\beta\alpha}L^{\beta}_{\phantom{\beta}\mu\nu}-L^{\beta}_{\phantom{\beta}\mu\alpha}L^{\rho}_{\phantom{\rho}\beta\nu}-L^{\beta}_{\phantom{\beta}\nu\alpha}L^{\rho}_{\phantom{\rho}\mu\beta}.$ We can then write $\Lambda=\Lambda(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})$ without loss of generality.
@@
    were each $\Lambda_{(k)}$ is a homogeneous function of degree $k$ in $\boldsymbol{g}$ and its derivatives; and the same property holds for $V$.
@@ V.3 Bootstrapping of $\Lambda$CDM
We want to construct here the Lagrangian density obtained by variational bootstrapping of (39). In the following, we will apply the variational bootstrapping procedure to the gravity+dark sector only. To include regular matter, one should just add the corresponding matter Lagrangian to the one obtained below. According to (44) and the above hypotheses, we thus need to compute
@@
and [^10] Note – see Appendix B – that under the rescaling $g_{\mu\nu}\mapsto ug_{\mu\nu}$, the Christoffel symbols $\accentset{\circ}{\Gamma}^{\rho}_{\mu\nu}$ remain invariant.
@@
Here again, according to hypothesis 3, we want to rewrite $V^{\mu}(\boldsymbol{g},\mathbf{L},\accentset{\circ}{\nabla}\mathbf{L})$ as
@@ V.4.1 Variation with respect to $\boldsymbol{g}$
where $\simeq$ means equality up to boundary terms. Let us now come to the variation of (54) with respect to the metric. To this purpose, we calculate separately the variation of each term $\Lambda_{(k)}$. Using relation (91) in Appendix B, we get:
@@
where, to shorten the notation, we have defined $\left(\partial\Lambda_{(k)}\right)^{\alpha\phantom{\rho}\sigma\beta}_{\phantom{\alpha}\rho}\coloneqq\frac{\partial\Lambda_{(k)}}{\partial\left(\accentset{\circ}{\nabla}_{\alpha}L^{\rho}_{\phantom{\rho}\sigma\beta}\right)}$. Using again the relations from Appendix B (in particular (92) and (94)), we get that [^11] Along the computation, it is also particularly useful to use the identity $A^{\mu\nu\alpha}\left(B_{\mu\nu\alpha}-B_{\alpha\mu\nu}+B_{\nu\alpha\mu}\right)=\left(A^{\mu\nu\alpha}-A^{\alpha\mu\nu}+A^{\nu\alpha\mu}\right)B_{\nu\alpha\mu}$ valid for any two objects with three indices.
$\delta_{g}\left(\Lambda_{(k)}\sqrt{-\lvert g\rvert}\right)\simeq\sqrt{-\lvert g\rvert}\ T^{\mu\nu}_{(k)}\ \delta g_{\mu\nu},$ (61)
with
$T_{\left(k\right)}^{\mu\nu}$ $=$ $\frac{\partial\Lambda_{\left(k\right)}}{\partial g_{\mu\nu}}+\frac{1}{2}\Lambda_{\left(k\right)}g^{\mu\nu}$
$+\frac{1}{2}\accentset{\circ}{\nabla}_{\alpha}\left\{\left(\partial\Lambda_{\left(k\right)}\right)_{~{}\rho}^{\tau~{}\sigma\beta}\left(-L_{~{}\sigma\beta}^{\gamma}g^{\rho\xi}B_{\xi\gamma\tau}^{\mu\nu\alpha}+L_{~{}~{}\beta}^{\rho\gamma}B_{\gamma\sigma\tau}^{\mu\nu\alpha}+L_{~{}\sigma}^{\rho~{}\gamma}B_{\gamma\beta\tau}^{\mu\nu\alpha}\right)\right\}$
where $B_{\xi\gamma\tau}^{\mu\nu\alpha}$ is a combination of Kronecker deltas:
$B_{\xi\gamma\tau}^{\mu\nu\alpha}:=\frac{1}{2}\left(\delta_{\xi}^{\mu}\delta_{\gamma}^{\nu}\delta_{\tau}^{\alpha}+\delta_{\xi}^{\nu}\delta_{\gamma}^{\mu}\delta_{\tau}^{\alpha}-\delta_{\tau}^{\mu}\delta_{\gamma}^{\nu}\delta_{\xi}^{\alpha}-\delta_{\tau}^{\nu}\delta_{\gamma}^{\mu}\delta_{\xi}^{\alpha}+\delta_{\xi}^{\mu}\delta_{\tau}^{\nu}\delta_{\gamma}^{\alpha}+\delta_{\xi}^{\nu}\delta_{\tau}^{\mu}\delta_{\gamma}^{\alpha}\right).$ (63)
@@
$\delta_{g}\left(V_{(kl)}\sqrt{-\lvert g\rvert}\right)\simeq\sqrt{-\lvert g\rvert}\ T^{\mu\nu}_{(kl)}\ \delta g_{\mu\nu},$ (65)
where the expression for $T^{\mu\nu}_{(kl)}$ is obtained by performing the replacement $\Lambda_{(k)}\to V_{(kl)}$ in (V.4.1).
Combining (59), (61) and (65), we get that the metric field equation for the Lagrangian (46), in the absence of matter source, is
@@
where $\boldsymbol{T}^{(k)}=T^{(k)}_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ and $\boldsymbol{T}^{(kl)}=T^{(kl)}_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ are as above.
@@ V.4.2 Variation with respect to $\mathbf{L}$
$\delta_{L}\left(\Lambda_{(k)}\sqrt{-\lvert g\rvert}\right)=\delta_{L}\Lambda_{(k)}\ \sqrt{-\lvert g\rvert}=\left(\frac{\partial\Lambda_{(k)}}{\partial L^{\rho}_{\phantom{\rho}\sigma\beta}}\delta L^{\rho}_{\phantom{\rho}\sigma\beta}+\left(\partial\Lambda_{(k)}\right)^{\alpha\phantom{\rho}\sigma\beta}_{\phantom{\alpha}\rho}\ \delta_{L}\left(\accentset{\circ}{\nabla}_{\alpha}L^{\rho}_{\phantom{\rho}\sigma\beta}\right)\right)\sqrt{-\lvert g\rvert}.$ (67)
Using identity (97) and integration by parts, we can write
$\delta_{L}\left(\Lambda_{(k)}\sqrt{-\lvert g\rvert}\right)\simeq\sqrt{-\lvert g\rvert}\ \left(\Psi_{(k)}\right)^{\phantom{\rho}\sigma\beta}_{\rho}\ \delta L^{\rho}_{\phantom{\rho}\sigma\beta},$ (68)